OpenStudy (anonymous):

Calculus 1!! find the intervals on which f is continuous: 1/(x+5^2)+10 Answer is: (-∞,∞) So I did this: 1/(x+5)² +10 --> (x² +10x+25) -->(5,5) But I am getting on the wrong track!! What steps do I take to get (-∞,∞)?? Can anyone help for a best answer? Please and thank you!!

OpenStudy (anonymous):

You must have a typo because the function you gave is not continuous on (-inf,inf)

OpenStudy (anonymous):

I know!!! The answer key said that it was supposed to be (-inf, inf). So what is the real answer supposed to be??? I got (-∞,-5),(-5,∞). But the other two answer choices were; b) (-∞,-40),(-40,∞) and d) (-∞,35),(35,∞)

OpenStudy (anonymous):

The plus 10 is supposed to be part of the denominator by the way! The problem is supposed to look like (1)/[(x+5²)+10]

OpenStudy (anonymous):

find the domian and that;s your answer

OpenStudy (unklerhaukus):

\[\frac1{(x+5^2)+10}\\ \frac1{(x+25)+10}\\ \frac1{x+35}\\ \] i think there is typo in question

OpenStudy (anonymous):

I think the question is:\[\frac{ 1 }{(x+5)^{2}+10 }\] If that is the case, \[(x+5)^{2}\] will never be negative. Therefore the min value of the denominator is 10. So for all values of x, the denominator will not be 0. Therefore it is continuous on \[(-\infty, \infty)\]